OpenStudy (anonymous):

5 questions for algebra please help!!!

OpenStudy (dayakar):

where are the questions one by one

OpenStudy (anonymous):

1.What is the slope of the line that passes through the pair points (-7/2,-3) and (-5,5/2) a.6/22 b.-6/22 c.22/6 d-22/6 2.tell whether the lines for each pair of equations are parallel,perpendicular,or niether Y=-3x+7 -2x+6y=3 a.parallel b.perpendiclar c.neither 3.tell whether the lines for each pair of equations are parallel,perpendicular,or neither Y=-1/4x+10 -2x+8y=6 a.parallel b.perpendicular c.neither 4. the table shows the number of miles driven over time. Time(hours)|distance(miles) 4 232 6 348 8 464 10 580 Express the relashionship between distance and time in simplified form as a unit rate. Determine which statement correctly interprets this relashionship. a. 232; your care travels 232 miles. b. 1/58; your car travels 58 miles in 1 hour. c. 58/1; your car travels 58 miles in 1 hour d. 1/10; your car travels for 10 hours 5. For the data in the table, does y vary directly with x? if it does write an equation for the direct variation. X | Y 10 12 15 18 20 24 a. Yes; y=1.2x b. Yes; y=2x c. Yes; y=x+1

OpenStudy (anonymous):

@dayakar

OpenStudy (michele_laino):

hint: the slope \(m\) of a line which passes at points: \[\left( {{x_1},{y_1}} \right),\quad \left( {{x_2},{y_2}} \right)\] is given by the subsequent formula: \[m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}}\]

OpenStudy (anonymous):

Yea i know the slope formula but when i did it i got none of the anwsers so im confused

OpenStudy (michele_laino):

let's suppose this: \[\left( {{x_1},{y_1}} \right) = \left( { - 5,5/2} \right),\quad \left( {{x_2},{y_2}} \right) = \left( { - 7/2, - 3} \right)\]

OpenStudy (anonymous):

so -7/2-(-5)/-3-5/2 right?

OpenStudy (michele_laino):

then after a simple substitution, we get: \[\huge \begin{gathered} m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{ - 3 - \frac{5}{2}}}{{ - \frac{7}{2} - \left( { - 5} \right)}} = \hfill \\ \hfill \\ = \frac{{\frac{{\left( {\left( { - 3} \right) \cdot 2} \right) - 5}}{2}}}{{\frac{{ - 7 + \left( {5 \cdot 2} \right)}}{2}}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

i got -5 1/2/1 1/2

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

such result can be rewritten as follows: \[\huge \begin{gathered} m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{ - 3 - \frac{5}{2}}}{{ - \frac{7}{2} - \left( { - 5} \right)}} = \hfill \\ \hfill \\ = \frac{{\frac{{\left( {\left( { - 3} \right) \cdot 2} \right) - 5}}{2}}}{{\frac{{ - 7 + \left( {5 \cdot 2} \right)}}{2}}} = \frac{{ - \frac{{11}}{2}}}{{\frac{3}{2}}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now, please multiply both numerator and denominator, of last side, by \(4\)

OpenStudy (anonymous):

it would be -22/6 right?

OpenStudy (michele_laino):

 that's right!

OpenStudy (anonymous):

Can you help me with number 2?

OpenStudy (michele_laino):

second question: two lines, for example: \(y=m_1x+q_1\) and \(y=m_2x+q_2\) are parallel if \(m_1=m_2\)

OpenStudy (anonymous):

so is number 2 perpendicular

OpenStudy (michele_laino):

we have this: the first line is \(y=-3x+7\), so the first slope is \(m_1=-3\) whereas the second line is \(y=(1/3)x+1/2\) so the second slope is \(m_2=1/3\)

OpenStudy (michele_laino):

now, please what is the product: \(m_1 \times m_2=...?\)

OpenStudy (anonymous):

m1=1/2 m2 m2=1/2 so they have the same y intercept correct?

OpenStudy (marieblossom16):

for #2 you have to put both equations in slope intercept for and if they have the same y intersept then it perp if it has the same slope its par.

OpenStudy (anonymous):

so im correct the second one is perpendicular because it has the same slope just reciprocal of each other and the y intercept is the same?

OpenStudy (michele_laino):

hint: we can write this: \[\huge {m_1} \times {m_2} = \frac{1}{3} \times \left( { - 3} \right) = ...?\]

OpenStudy (anonymous):

-1

OpenStudy (michele_laino):

that's right they are perpendicular. Nevertheless the y-intercept of the first line is 7, whereas the y-intercept of the second line is 1/2

OpenStudy (anonymous):

and im guessing number 3 is paralell

OpenStudy (anonymous):

i think number 3 is parallel because the slopes are the same but the y intercepts are different am i wrong?

OpenStudy (michele_laino):

we can rewrite the second equation as below: \(y=(1/4)x+3/4\) so the slope of the second line is \(1/4\), whereas the slope of the first line is \(-1/4\)

OpenStudy (michele_laino):

now we can see that: \(m_1\) is different from \(m_2\), furthermore \(m_1 \times m_2=-1/16\), so what can you conclude?

OpenStudy (anonymous):

paralell

OpenStudy (michele_laino):

are you sure? parallel lines have the same slope

OpenStudy (anonymous):

ohh so it would be neither

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

i didnt see the negative sign for the second slope ;)

OpenStudy (michele_laino):

ok! :)

OpenStudy (michele_laino):

now, for the fourth exercise, we can see that the ratio: \[\frac{{{\text{distance}}}}{{{\text{time}}}}\] is constant for each ordered pair. Please what is the value of such constant?

OpenStudy (michele_laino):

namely, please compute these ratios: \[\Large \begin{gathered} \frac{{232}}{4} = ...? \hfill \\ \hfill \\ \frac{{348}}{6} = ...? \hfill \\ \hfill \\ \frac{{464}}{8} = ...? \hfill \\ \hfill \\ \frac{{580}}{{10}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

they all equal 58

OpenStudy (anonymous):

so the anwser is.... 58/1 or 1/58 i got confused on that

OpenStudy (michele_laino):

yes! correct! so the requested relation is: \[\large \frac{{{\text{distance}}}}{{{\text{time}}}} = 58\] or: \[\large {\text{distance}} = 58 \times {\text{time}}\]

OpenStudy (michele_laino):

now, if \(time = 1\) hour, what is the \(distance\)? \[\large {\text{distance}} = 58 \times {\text{time}} = 58 \times 1 = ...?\]

OpenStudy (anonymous):

58/1

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

for 5 im really confused

OpenStudy (anonymous):

im not good at direct variations

OpenStudy (anonymous):

5. For the data in the table, does y vary directly with x? if it does write an equation for the direct variation. X | Y 10 12 15 18 20 24 a. Yes; y=1.2x b. Yes; y=2x c. Yes; y=x+1 d. No; ydoes not vary directly with x.

OpenStudy (michele_laino):

hint: again please compute these ratios: \[\huge \begin{gathered} \frac{{12}}{{10}} = ...? \hfill \\ \hfill \\ \frac{{18}}{{15}} = ...? \hfill \\ \hfill \\ \frac{{24}}{{20}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

oooooh 1.2 so ye=1.2x

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thanks ;) you are a really good helper

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

can i ask you another question? @Michele_Laino

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

For the data in the table does y vary directly with x? if it does, write an equation for the direct variation X | Y 40 32 28 16 16 12 a. Yes; y=2x b. Yes; y= 0.5x c. Yes’ y=1.5x d. No; y does not vary directly with x

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

Desribe how the graphs of y=|x| and y=|x+5| are related a. The graphs have the same shape. The y intercept of y=|x| is 0 and the y intercept of the second graph is 5 b. The graphs have the same shape. The y intercept of y =|x| is 0 and the y intercept of the second graph is -5 c. The two graphs are the same. d. The graphs have the same shape. The y intercept of y=|x| is 0 and the x intercept if the second graph is 5

OpenStudy (anonymous):

Desribe how the graphs of y=|x| and y=|x|-5 are related a. The graphs have the same shape. The y intercept of y=|x| is 0 and the y intercept of the second function is 5 b. the functions have the same y intercept. The second function is steeper tha y=|x| c. The two functions are the same. d. The graphs have the same shape. The y intercept of y=|x| is 0 and the x intercept if the second function is - 5

OpenStudy (anonymous):

For the first one, i would suggest multiplying your x's in to your equations.

OpenStudy (anonymous):

i think it is D

OpenStudy (anonymous):

it doesnt vary directly with x @Lovelarap

OpenStudy (anonymous):

yup

OpenStudy (michele_laino):

as we can see the y-coordinate of the graph \(y=|x|-5\) is 5 points less than the corresponding graph \(y=|x|\)

OpenStudy (anonymous):

for the last two, try this document:

OpenStudy (anonymous):

so my anwser is d. The graphs have the same shape. The y intercept of y=|x| is 0 and the x intercept if the second function is - 5

OpenStudy (michele_laino):

that's right! I think that in your textbook there is a typo, since the correct statement is: The graphs have the same shape. The y intercept of y=|x| is 0 and the \(y\) intercept if the second function is - 5

OpenStudy (anonymous):

oh i wrote it wrong it said y intercept

OpenStudy (michele_laino):

ok! :) it is option d

OpenStudy (anonymous):

and for Desribe how the graphs of y=|x| and y=|x|-5 are related a. The graphs have the same shape. The y intercept of y=|x| is 0 and the y intercept of the second graph is 5 b. the functions have the same y intercept. The second functions is steeper tha y|x| c. The two graphs are the same. d. The graphs have the same shape. The y intercept of y=|x| is 0 and the x intercept if the second graph is 5 my anwser is either a or d but im not sure

OpenStudy (anonymous):

oops i ment y=|x| and y=|x+5| are related

OpenStudy (anonymous):

a. The graphs have the same shape. The y intercept of y=|x| is 0 and the y intercept of the second graph is 5 b. The graphs have the same shape. The y intercept of y =|x| is 0 and the y intercept of the second graph is -5 c. the two graphs are the same. d. The two graphs are the same. d. The graphs have the same shape. The y intercept of y=|x| is 0 and the x intercept if the second graph is 5

OpenStudy (michele_laino):

in order to find the y.intercept, we have to replace \(x=0\). So the y-intercept of the first graph,is: \(y=|x|=|0|=...?\) whereas the y-intercept of the second graph is: \(y=|x+5|=|0+5|=...?\)